I need help with quadratic word problems.
Noah is trying to qualify for a race. A driver needs to complete 8 laps of a seven mile track at an average speed of 105.66 mph to qualify. Noah has already completed the first three laps. His average speed so far is 92 mph. What must his average speed be for the rest of the race to qualify?

Answer by stanbon(75887)

A driver needs to complete 8 laps of a seven mile track at an average speed of 105.66 mph to qualify. Noah has already completed the first three laps. His average speed so far is 92 mph. What must his average speed be for the rest of the race to qualify?
--------------------------------------
He has gone 21 miles at 92 mph.
His elapsed time is 21/92 hr.
------------------------------------
The 8 lap distance is 56 miles.
To average 105.66 means total time must be 56/105.66 hr.
------------------------------------
Time remaining after 3 laps is (56/105.66)-(21/92) = 0.30174 hrs.
----------------------------------------
Distance remaining after 3 laps is 5*7 = 35 miles.
-------------------------------------------
Required rate for the remaining 35 miles is 35/0.30174 = 115.994 mph
===============================================================
Cheers,
Stan H.